How much heat (in kJ) is released when 15.0 L of CO at 85°C and 112 kPa reacts with 18.5 L of H2 at 75°C and 744 torr?


1
Expert's answer
2021-11-17T15:27:03-0500

"CO+2H_2\to CH_3OH"

"V_{CO}=15L=0.015m^3"

"P=112000Pa"

"T=85^0C=358,15K"

"PV = nR T"

"n=(PV)\/(RT)"

"n= (1120000,015)\/(8,314358,15)"

n(CO)=0,564mol

"V_{CO}=18,5L=0,0185m^3"

"P=744torr=99191,84Pa"

"T=75^0C =348,15K"

"PV =n RT"

"n=(PV)\/(RT)"

"n=(99191,840,0185)\/(8,314348,15)"

"n(H_2)=0,634mol"

It can be seen that "H_2" is the limiting reactant

"n(CH_3OH)=1\/2n(H_2)=1\/20,634MOL=0,317mol"

"\Delta H_{rxn}=\Delta H_f(CH_3OH)-\Delta H_f(CO)=-238,6-(-110,5)=-128,1kJ/mol"

"E=\Delta H_{rxn}\times n =-40,6kJ"